OpenStudy (anonymous):

solve the system of linear equations bellow: show work y=-2x+19 y=x+7

OpenStudy (blurbendy):

you know y = x + 7, so put x + 7 in for y in the top equation and solve for x

OpenStudy (anonymous):

so it would look like x+7=2x+19

OpenStudy (blurbendy):

-2x not +2x

OpenStudy (anonymous):

so x would be -12 how do i find y

OpenStudy (blurbendy):

now you know x = -12, so you -12 in for x in the bottom equation and solve for y

OpenStudy (anonymous):

so it would look like y=-12+7

OpenStudy (blurbendy):

yes

OpenStudy (anonymous):

thanks for the help:)

OpenStudy (blurbendy):

yw